# A Self-Supervised Model for Multi-modal Stroke Risk Prediction

### v1

@@ 3.1.1 Supervised learning methods
To benchmark our proposed method, we implement two state-of-the-art, supervised image-based models, namely ResNet50 He et al. [2020] and DenseNet121 Huang et al. [2017], two supervised tabular data approaches, namely a two-layer tabular MLP model and a tabular transformer encoder inspired from Du et. al. (2024) Du et al. [2024]. We conduct an ablation study using a supervised MLP model with various feature combinations to identify the optimal feature set. This process helps us to select the final combination of features for improved model performance. The combinations include: $i)$ clinical tabular data only, spanning the previously mentioned categories (clinical), $ii)$ clinical data with brain extracted IDPs (clinical + brain IDPs), and $iii)$ clinical data, brain IDPs and lesion IDPs (clinical + brain IDPs + lesion IDPs). Furthermore, we implement three supervised, multi-modal (imaging-tabular) learning models, namely a simple concatenation fusion model (CF) Spasov et al. [2019], a CF model integrated with the tabular transformer encoder inspired by the work from Du et. al. (2024) Du et al. [2024] (CF + Transformer), and DAFT model Wolf et al. [2022]. All models employing an imaging encoder are implemented with ResNet50 as backbone. DAFT block is integrated within ResNet50 from the third stage onwards. To alleviate over-fitting, an early stopping strategy is adopted, with a minimal delta (divergence threshold) of $1\times 10^{-4}$, a maximal number of epochs of 50, and a patience of 15 epochs.
@@ 3.1.2 Self-supervised learning methods
Our model is compared against leading, self-supervised contrastive solutions, including: $i)$ the unimodal, image-based SimCLRChen et al. [2020] approach, $ii)$ the unimodal, tabular data-based SCARF Bahri et al. [2022] approach, and $iii)$ the multi-modal, CLIP method (without ITM loss). The hyperparameters and training configurations for all SSL pre-training approaches are adapted for our specific dataset and task, and are obtained through hyper-parameter search. All models are pre-trained for 100 epochs using an Adam optimizer Kingma and Ba [2015]. The learning rate is warmed up linearly for 10 epochs and decayed following a cosine annealing scheduler. For all methods, as in the CLIP+ITM method, the image augmentation rate is 95% and the tabular corruption rate 0.3 during pre-training and 80% and 0.3 during downstream fine tuning. SimCLR is trained using the NTXent objective Chen et al. [2020] and the temperature parameter is kept to 0.1. Hidden and projected dimensions are respectively 2048 and 128 for both modalities Chen et al. [2020]. The same parameters are used for SCARF Bahri et al. [2022]. Learning rates are chosen with a sweep through a range of learning rates, by tracking the validation loss. Weight decay and dropout rate are added depending on the level of overfitting observed at the validation loss. The same early stopping strategy is employed as in the supervised learning methods. The downstream fine-tuning is using the same parameters as the presented method, using only a single modality classifier for unimodal SSL pre-trained methods and a fused representation vector with a single linear classifier for the CLIP only model. Trainable models in each SSL unimodal method means that the other modality is incorporated during fine-tuning as a full trainable model. The employed batch size for all methods is 6. All SSL models are pretrained on a Tesla V100-SXM2 (32GB, 42 CPUs), and inference is performed on an NVIDIA GeForce RTX 4090 (24GB, 62 CPUs). Pretraining took $\sim$ 24 hours, while fine-tuning took less than 1 hour.
@@ 3.2 Interpretability and qualitative analysis
Embeddings visualization is done using a two-dimension Uniform Manifold Approximation and Projection (UMAP) technique McInnes et al. [2018], to evaluate the quality of the generated latent space embedding after pre-training, using validation samples. Such approach allows to qualitatively assess the latent space representation and data distribution after encoding each modality with either a unimodal or multi-modal pre-trained model, giving some hints about the successfullness of the learning strategies. A latent space embedding size of 2048 dimensions is produced.
For qualitative analysis, we generated imaging heatmaps using the Gradient-weighted Class Activation Mapping (GradCAM) technique Selvaraju et al. [2017] to visualize the regions in each slice that contributed most significantly to the model’s predictions across given brain MRI volume. GradCAM Selvaraju et al. [2017] heatmaps are normalized in the range of 0 to 1 and are upsampled with trilinear interpolation to match the original image space. The 7th layer of the ResNet50 encoder is used to allow capturing high-level features and spatial structure that is suitable for visualization. The most informative slice (defined as the one accounting with the highest heatmap activation scores) for each view in axial, sagittal, or coronal plane is generated. We use the 3D GradCAM implementation from MONAI.
@@ 3.3 Performance assessment
All models are evaluated through the area under the Receiver Operating Characteristic (ROC) curve. Binary classification metrics, namely balanced accuracy, F1-Score, and sentitivty, are included. Classification metrics are reported at the Youden-index operating point ($J=\text{Sensitivity}+\text{Specificity}-1$) retrieved from the (validation set) ROC curve. The metrics are chosen bearing in mind that potential clinical applications of this study could serve as screening and risk stratification tools, where the models sensitivity plays an important role to avoid missing positive stroke cases.
@@ 4.1 Benchmarking
Table 1: Feature selection with supervised MLP results using different combinations of tabular data. ROC-AUC: area under the receiver operating characteristic curve; bAcc (%): balanced Accuracy; Se (%): sensitivity. For each metric, the best-performing method is highlighted in bold and the second-best is underlined.
@@
When comparing SSL strategies, it is evident that fine-tuning both data modalities in multi-modal approaches significantly boosts performance. The performance gap is considerable when comparing these multi-modal models with unimodal image-based models, showing that image data alone is insufficient for effectively addressing the task. The best performing method is the CLIP+ITM model, performing better than all unimodal (tabular and imaging) SSL methods. Interestingly, DAFT performs similar to the multi-modal SSL methods in terms of ROC-AUC and balanced accuracy, although exhibits poor F1-score and sensitivity results. There is no clear difference in performance between trainable and frozen settings across all models. We hypothesize this is because the pre-trained models have already developed robust, transferable representations, making fine-tuning less impactful. Additionally, the small size of the fine-tuning dataset may limit the effectiveness of further learning beyond what was achieved during pre-training. Besides, it could be hypothesized that freezing the model may serve as a form of regularization, helping to mitigate overfitting, particularly in this setting with limited labeled data.
Table 2: Benchmarking performance results. F and T denote frozen and trainable pre-trained encoders. ROC-AUC: area under the receiver operating characteristic curve; bAcc (%): balanced Accuracy; Se (%): Sensitivity. For each metric, the best-performing method is highlighted in bold and the second-best is underlined. The overall best performing method is highlighted in gray.
Model Tabular Image Metrics (%) AUC bAcc F1 Se (a) Supervised Image ResNet-50 He et al. [2016] - T 63.25 57.08 60.01 65.22 DenseNet121 Huang et al. [2017] - T 66.79 66.79 69.90 78.26 (b) Supervised Tabular MLP T - 71.37 60.31 65.96 67.39 Transformer Du et al. [2024] T - 64.38 62.21 47.62 32.61 (c) Supervised multi-modal Concat Fuse (CF) Spasov et al. [2019] T T 65.26 60.29 62.63 67.39 Concat Fuse (CF) [w/ Transformer] Spasov et al. [2019], Du et al. [2024] T T 63.48 52.08 66.17 95.65 DAFT Wolf et al. [2022] T T 73.82 63.51 69.57 65.01 (d) SSL Image SimCLR Chen et al. [2020] - F 64.99 52.38 33.33 28.91 SimCLR Chen et al. [2020] - T 65.59 55.67 43.83 34.78 SimCLR Chen et al. [2020] T F 72.02 65.56 64.44 63.04 SimCLR Chen et al. [2020] T T 72.11 65.56 64.44 63.04 (e) SSL Tabular SCARF Bahri et al. [2022] F - 71.18 62.42 63.91 67.39 SCARF Bahri et al. [2022] T - 70.35 64.48 62.92 60.87 SCARF Bahri et al. [2022] F T 72.16 62.16 43.48 53.34 SCARF Bahri et al. [2022] T T 72.02 67.85 73.08 78.26 (f) SSL multi-modal CLIP Hager et al. [2023] T T 73.41 61.5 67.24 80.78 CLIP Hager et al. [2023] F F 73.54 71.00 70.97 71.74 CLIP+ITM Du et al. [2024], Li et al. [2021] T T 74.42 71.11 74.22 84.78 CLIP+ITM Du et al. [2024], Li et al. [2021] F F 74.75 62.77 67.29 76.60
@@ 4.2.1 Embeddings visualization
Figure 2 shows the UMAP embeddings distribution for unimodal and multi-modal data models. On one hand, it can be observed that in Fig. 2 A, there is a clear distinction between (unimodal learnt) tabular and imaging data modalities, with data samples clustered by data-type. In this case, the embeddings generated from imaging data and tabular data are significantly different from each other in the feature space when generated with a unimodal pre-trained model (i.e., either SCARF or SimCLR). The tight clustering of red points suggests that the tabular data embeddings are more homogeneous and possibly more concentrated in the feature space compared to the broad representation of brain MRI images. Therefore, unimodal-data encoders have learned modality-specific features, without capturing interactions between them. On the other hand, in Fig. 2 B the UMAP plot obtained for the best performing multi-modal model (CLIP+ITM) is shown. In this case, there is significant overlap between the tabular and imaging embeddings, suggesting that the model has found common representations for the two different data types, either via shared visual features or via learning associated clinical patterns in tabular and brain MRIs. Thus, CLIP+ITM is able to encode the underlying patient representation in a common latent space by reducing data augmentation noise. Still, there are data-points in the plot having distinct representations within each modality, suggesting that the model could not project them to the modality-shared latent space. The broad distribution of points across the entire UMAP space suggests that the embeddings capture a wide variety of features from both imaging and tabular data, rather than collapsing all data points into a narrow cluster. These results expose the enhanced performance of the multi-modal SSL strategy by projecting diverse data modalities into a shared embedding space, and thus suggesting a better model starting point for downstream analysis.
@@
Figure 2: 2D UMAP projections of tabular and imaging embeddings from the validation set, using (A) unimodal pre-trained tabular and imaging encoders and (B) multi-modal pre-trained tabular and imaging encoders.
4.2.2 Imaging heatmaps

### v2

@@ 3.1.1 Supervised learning methods
To benchmark our proposed method, we implement two state-of-the-art, supervised image-based models, namely ResNet50 [He2020] and DenseNet121 [Huang2017], two supervised tabular data approaches, namely a tabular MLP model and a tabular transformer encoder inspired by Du et al. (2024) [Du2024]. We conduct an ablation study using a supervised MLP model with various feature combinations to identify the optimal feature set. This process helps us to select the final combination of features for improved model performance. The combinations include: $i)$ clinical records only, spanning the previously mentioned categories (referred to as "clinical"), $ii)$ clinical records with brain extracted IDPs ("clinical + brain IDPs"), and $iii)$ clinical records, brain IDPs and lesion IDPs ("clinical + brain IDPs + lesion IDPs"). Furthermore, we implement three supervised, multi-modal (imaging-tabular) learning models, namely a simple concatenation fusion model (CF) [Spasov2019], a CF model integrated with the tabular transformer encoder inspired by the work from Du et al. (2024) [Du2024] (CF + Transformer), and DAFT model [Wolf2022]. All models employing an imaging encoder are implemented with ResNet50 as a backbone. The DAFT block is integrated within ResNet50 from the third stage onward. To alleviate overfitting, an early stopping strategy is adopted, with a minimal delta (divergence threshold) of $1\times 10^{-4}$, a maximal number of epochs of 50, and a patience of 15 epochs.
@@ 3.1.2 Self-supervised learning methods
Our model is compared against leading, self-supervised contrastive solutions, including: $i)$ the uni-modal, image-based SimCLR[Chen2020] approach, $ii)$ the uni-modal, tabular data-based SCARF [Bahri2022] approach, and $iii)$ the multi-modal, CLIP method (without ITM loss). The hyperparameters and training configurations for all SSL pre-training approaches are adapted for our specific dataset and task, and are obtained through hyper-parameter search. All models are pre-trained for 100 epochs using an Adam optimizer [Kingma2015]. The learning rate is warmed up linearly for 10 epochs and decayed following a cosine annealing scheduler. For all methods, the image augmentation rate is 95% and the tabular corruption rate is 0.3 during pre-training and 80% and 0.3 during supervised fine tuning. SimCLR is trained using the NTXent objective [Chen2020] and the temperature parameter is kept to 0.1. Hidden and projected dimensions are respectively 2048 and 128 for both modalities [Chen2020]. The same parameters are used for SCARF [Bahri2022]. Learning rates are chosen with a sweep through a range of learning rates, by tracking the validation loss. Weight decay and dropout rate are added depending on the level of overfitting observed at the validation loss. The same early stopping strategy is employed as in the supervised learning methods. The downstream fine-tuning employs the same parameters as the pre-training method, using a single modality classifier for the unimodal SSL pre-trained methods and a fused representation vector combined with a single linear classifier for the CLIP-only model. Trainable models in each SSL uni-modal method means that the other modality is incorporated during fine-tuning as a full trainable model. The employed batch size for all methods is 6. All SSL models are pretrained on a Tesla V100-SXM2 (32GB, 42 CPUs), and inference is performed on an NVIDIA GeForce RTX 4090 (24GB, 62 CPUs). The pretraining lasts around $\sim$ 24 hours, while fine-tuning less than 1 hour. Tables detailing the training parameters used during both pre-training and fine-tuning are included as Tables S4 and S5 in the Supplementary Material.
@@ 3.2 Interpretability and qualitative analysis
Embeddings visualization is done using a two-dimension Uniform Manifold Approximation and Projection (UMAP) technique [McInnes2020], to evaluate the quality of the generated latent space embedding after pre-training, using validation samples. Such approach allows to qualitatively assess the latent space representation and data distribution after encoding each modality with either a uni-modal or multi-modal pre-trained model, providing insights into the effectiveness of the learning strategies. A latent space embedding size of 2048 dimensions is produced.
For the qualitative analysis, we are interpreting the model’s predictions with respect to both MRI and clinical features modality. First, to analyze the importance of the individual tabular features in generating the embeddings, we compute the integrated gradients over the test samples with a multi-modal pre-trained tabular encoder, aggregating the importance of each feature across all embeddings dimensions, specifically handling categorical features by summing the mean importance over all choices. Integrated gradients are computed with the package Captum integrated with pytorch [kokhlikyan2020captum]. Then, we generate imaging heatmaps using the Gradient-weighted Class Activation Mapping (GradCAM) technique [Selvaraju2017] to visualize the regions in each slice that contributed most significantly to the model’s predictions across the brain MRI volumes. GradCAM [Selvaraju2017] heatmaps are normalized in the range of 0 to 1 and are upsampled with trilinear interpolation to match the original image space. The 7th layer of the ResNet50 encoder is used to allow capturing high-level features and spatial structure that is suitable for visualization. The most informative slice (defined as the slice with the highest heatmap activation scores) for each view in axial, sagittal, or coronal plane is generated. We use the 3D GradCAM implementation from MONAI [Cardoso2022]. Pursuing our analysis, we generate boxplots comparing the normalized mean GradCAM activations for healthy versus stroke individuals on WMH-segmented brain MRIs and on whole T2-FLAIR brain MRI scans.
@@ 3.3 Performance assessment
All models are evaluated through the area under the Receiver Operating Characteristic (ROC) curve. Binary classification metrics, namely (balanced) accuracy, F1-Score, and sensitivity, are included. Classification metrics are reported at the Youden-index operating point ($J=\text{Sensitivity}+\text{Specificity}-1$) retrieved from the (validation set) ROC curve. The metrics are chosen bearing in mind that potential clinical applications of this study could serve as screening and risk stratification tools, where the model’s sensitivity plays an important role to avoid missing positive stroke cases.
@@ 4.1 Benchmarking
Table 3: Feature selection with supervised MLP using different combinations of tabular data. ROC-AUC: area under the receiver operating characteristic curve; Acc (%): Accuracy; Se (%): sensitivity. For each metric, the best-performing method is highlighted in bold and the second-best is underlined.
@@
Table 4: Benchmarking performance results. F and T denote frozen and trainable pre-trained encoders, respectively. ROC-AUC: area under the receiver operating characteristic curve; Acc (%): Accuracy; Se (%): Sensitivity. For each metric, the best-performing method is highlighted in bold and the second-best is underlined. The overall best performing method is highlighted in gray.
Model Tabular Image Metrics (%) AUC Acc F1 Se (a) Supervised Image ResNet-50 [He2016] - T 63.25 57.08 60.01 65.22 DenseNet121 [Huang2017] - T 66.79 66.79 69.90 78.26 (b) Supervised Tabular MLP T - 71.37 60.31 65.96 67.39 Transformer [Du2024] T - 64.38 62.21 47.62 32.61 (c) Supervised multi-modal Concat Fuse (CF) [Spasov2019] T T 65.26 60.29 62.63 67.39 Concat Fuse (CF) [w/ Transformer] T T 63.48 52.08 66.17 95.65 DAFT [Wolf2022] T T 73.82 63.51 69.57 65.01 (d) SSL Image SimCLR [Chen2020] - F 64.99 52.38 33.33 28.91 SimCLR [Chen2020] - T 65.59 55.67 43.83 34.78 SimCLR [Chen2020] T F 72.02 65.56 64.44 63.04 SimCLR [Chen2020] T T 72.11 65.56 64.44 63.04 (e) SSL Tabular SCARF [Bahri2022] F - 71.18 62.42 63.91 67.39 SCARF [Bahri2022] T - 70.35 64.48 62.92 60.87 SCARF [Bahri2022] F T 72.16 62.16 43.48 53.34 SCARF [Bahri2022] T T 72.02 67.85 73.08 78.26 (f) SSL multi-modal CLIP [Hager2023] T T 73.41 61.5 67.24 80.78 CLIP [Hager2023] F F 73.54 71.00 70.97 71.74 Ours T T 74.42 71.11 74.22 84.78 Ours F F 74.75 62.77 67.29 76.60
@@
When comparing SSL strategies, it is evident that fine-tuning both data modalities in multi-modal approaches significantly boosts performance. The performance gap is considerable when comparing these multi-modal models with uni-modal image-based models, showing that image data alone is insufficient for effectively addressing the task. Our model is performing better than all uni-modal (tabular and imaging) SSL methods, as well as the method using only CLIP loss. Interestingly, DAFT performs similarly to the multi-modal SSL methods in terms of ROC-AUC and accuracy, although exhibiting poor F1-score and sensitivity results. There is no clear difference in performance between trainable and frozen settings across all models. We hypothesize this is because the pre-trained models have already developed robust, transferable representations, making fine-tuning less impactful. Additionally, the small size of the fine-tuning dataset may limit the effectiveness of further learning beyond what was achieved during pre-training. Besides, it could be hypothesized that freezing the model may serve as a form of regularization, helping to mitigate overfitting, particularly in this setting with limited labeled data.
To assess statistical significance of our model’s performance superiority on the true stroke prevalence from the original pre-training stroke distribution, we have performed Wilcoxon signed-rank tests comparing each model against ours in Table 5.The results indicate that our model significantly outperforms all other uni-modal and multi-modal unsupervised baselines, with $p$-values $<~{}0.01$.
Table 5: Test set performance of unsupervised models for stroke classification. Results were obtained via bootstrapping with replacement (1,000 iterations), preserving the real-world stroke prevalence in the UK Biobank ( 0.52 %). Statistical significance was assessed using the Wilcoxon signed-rank test, comparing the baseline models against our proposed approach. AUC: area under the receiver operating characteristic curve. BA: balanced accuracy. CI: confidence intervals. * denotes $p$-values $<~{}0.01$.
Model AUC [95% CI] BA [95% CI] F1-score [95% CI] Sensitivity [95% CI]
SimCLR 0.720 [0.667–0.772]* 0.593 [0.556–0.628]* 0.6456 [0.6213–0.7362]* 0.631 [0.530–0.730]*
SCARF 0.727 [0.672–0.780]* 0.615 [0.571–0.659]* 0.7108 [0.7024–0.7229]* 0.740 [0.650–0.830]*
CLIP 0.734 [0.680–0.787]* 0.656 [0.606–0.705]* 0.7127 [0.6983–0.7342]* 0.782 [0.700–0.860]*
Ours 0.748 [0.696–0.798] 0.721 [0.677–0.761] 0.7443 [0.7233–0.7652] 0.846 [0.770–0.910]
In terms of model performance with respect to the stroke onset time variable, shown in Figure 3(a), the highest accuracy is achieved for people developing stroke after 2-3 years. First, strokes occurring closer to the scan acquisition date might be driven by acute mechanisms, such as embolism from an undiagnosed cardiac condition or sudden plaque rupture [Adams1993]. These acute events may not leave detectable brain MRI changes prior to stroke onset. Second, the model might be biased towards more chronic changes, long-term stroke risk factors as the dataset is skewed towards longer timeframes. However, the small sample size (N=93), might also be insufficient to draw such conclusions. Further, in Figure 3(b), the high proportion of correctly predicted strokes in older age groups (particularly 70-80), as well as the higher F1 score suggests that age is not only the strongest predictor variable, but it may also correlate with features that the model recognizes more confidently as indicative of stroke. In the healthy group, this trend is less pronounced, indicating that the model’s outputs are less influenced by the individuals’ ages. Given the inherent relationship between age and stroke risk, we conducted a logistic regression analysis to ensure the model is not undesirably inferring age as a proxy for stroke risk rather than predicting an individual’s true stroke risk. This analysis included various combinations of predictor variables, such as patient age, total WMH volume in brain regions (a known mediator of the relationship between age and stroke risk), and the predicted probabilities from the best-performing model from Table 4. As shown in Table S3 in the supplementary materials, incorporating the predicted probabilities into the logistic regression model yields approximately a 20% improvement compared to models based solely on age or WMH volume. Age and WMH volume are complementary predictors: when combined, they perform slightly better than when used individually. However, their predictive capabilities remain modest. While these features provide valuable information, they do not achieve the discriminative power of the model’s predicted probabilities, which largely encapsulate the relevant predictive information. This finding suggests that additional factors - or more complex relationships captured by the model generating the predicted probabilities - play a more significant role in determining outcomes. In any case, it is important to note that the test set comprises 93 patients, of whom 64 are over 60 years old and 29 are under 60 years old. This distribution should be considered when interpreting the results.
(a) Model performance vs stroke onset time.
(b) Model performance vs age.
Figure 3: Distribution of metrics with respect to group partition. (A) Model accuracy with respect to stroke onset time for stroke patients. Number of patients in each group are given on top of bars. Stroke onset time is the time from scan acquisition to stroke onset. Note: intervals are left-inclusive and right-exclusive. (B) Distribution of the predictions with respect to the age group. Above the bars, the number of patients in each group is given and in the middle of each bar, the F1 score.
@@ 4.2.1 Embeddings visualization
Figure 4 shows the UMAP embeddings distribution for uni-modal and multi-modal data models. On the one hand, it can be observed that in Fig. 4 A, there is a clear distinction between (uni-modal learnt) tabular and imaging data modalities, with data samples clustered by data-type. In this case, the embeddings generated from imaging data and tabular data are different from each other in the feature space when generated with a uni-modal pre-trained model (i.e., either SCARF or SimCLR). The tight clustering of red points suggests that the tabular data embeddings are more homogeneous and possibly more concentrated in the feature space compared to the broad representation of brain MRI images. Therefore, uni-modal data encoders have learned modality-specific features, without capturing interactions between them. On the other hand, in Fig. 4 B the UMAP plot obtained using our method (multi-modal, using CLIP and ITM). In this case, there is a consequent overlap between the tabular and imaging embeddings, suggesting that the model has found common representations for the two different data types, either via shared visual features or via learning associated clinical patterns in tabular and brain MRIs. Thus, our method is able to encode the underlying patient representation in a common latent space by reducing data augmentation noise. Still, there are data-points in the plot having distinct representations within each modality, suggesting that the model could not project them to the modality-shared latent space. The broad distribution of points across the entire UMAP space suggests that the embeddings capture a wide variety of features from both imaging and tabular data, rather than collapsing all data points into a narrow cluster. These results expose the enhanced performance of the multi-modal SSL strategy by projecting diverse data modalities into a shared embedding space, and thus suggesting a better model starting point for downstream analysis.
@@
Figure 4: UMAP projections of tabular and imaging embeddings from validation set, using (A) uni-modal pre-trained tabular and imaging encoders and (B) multi-modal pre-trained tabular and imaging encoders.
4.2.2 Imaging heatmaps and tabular feature importance
The analysis of tabular features importance, shown in Figure 5(a) demonstrates that age and image-derived features, such as PWMHs and DWMHs, contribute significantly to the predictions in the multi-modal model, likely indicating their robust representation of brain pathology. Clinical features like physical activity, cholesterol level, blood pressure lowering medication, hypertension, systolic pressure, previous symptoms of atrial fibrillation and sex also play key roles, though their relative importance is lower compared to the imaging-derived features. The use of integrated gradients allows for a nuanced understanding of the contributions from both modalities, highlighting how brain MRI features complement clinical records in predicting outcomes. This balanced contribution supports the potential of multi-modal models to integrate complex, heterogeneous data sources effectively for robust clinical insights. However, unlike the variable "Age", showing a statistically significant difference between groups in LABEL:tab:demographics, and is a key contributor to the model’s decision-making (Figure 5(a)), the variable "BMI", although also statistically significant, is not among the most influential tabular features (Figure 5(a)). The observed differences could still introduce subtle biases if the model were to generalize to populations with different BMI distributions. We acknowledge this as a potential limitation and suggest that future work could explore domain adaptation or covariate adjustment techniques to further mitigate residual confounding [Austin2011PropensityScore].